OpenStudy (anonymous):

What does it mean to refer to a 20 in TVor 25 in-TV set. A 20 in TV set also has a width of 16 inches, what is the height

OpenStudy (anonymous):

TV are measured in diagonal

OpenStudy (anonymous):

20in is the diagonal. so use the pythagorean theorem (a^2 + b^2 = c^2) to solve.

OpenStudy (anonymous):

what is the height

OpenStudy (anonymous):

solve using pythagorean formula. width squared + height squared = diagonal squared. (hint: this looks like a 3-4-5 triangle).

OpenStudy (anonymous):

I came up with 17 inches

OpenStudy (anonymous):

that is not possible.

OpenStudy (anonymous):

\[20 = \sqrt{16^2 + height^2}\]\[400 = 16^2 + height^2\]\[400 = 256 + height^2\]\[144 = height^2\]\[height = ???\]

OpenStudy (anonymous):

the way I wrote it was 16^2+b^2=18^2, 256+b^2=324, b^2=324-256,256/2=68/2=14/2=17

OpenStudy (anonymous):

where'd you get 18? your question said the diagonal was 20.

OpenStudy (anonymous):

I think I added 16+2

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

I am confuse Ithis is my 1 time doing algebra

OpenStudy (anonymous):

OK. my first post may have confused you. I meant: \[a^2 + b^2 = c^2\](otherwise known as the pythagorean theorem)

OpenStudy (anonymous):

so \[16^2 + b^2 = 20^2\]

OpenStudy (anonymous):

but if you recognize that one side is 16, and the diagonal is 20, you might remember that this is a common 3-4-5 right triangle with sides that are 4x longer...

OpenStudy (anonymous):

so will it make it 4

OpenStudy (anonymous):

so you're telling me \[16^2 + 4^2 = 20^2\]???

OpenStudy (anonymous):

I still can't get it sorry

OpenStudy (anonymous):

\[a^2+b^2=c^2\]\[b^2=c^2-a^2\]\[b^2=20^2-16^2\]\[b^2=400-256\]\[b^2=144\]\[b=???\]

OpenStudy (anonymous):

you understand that \[a^2=a \times a\]right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it 72

OpenStudy (anonymous):

Then I divide by 2

OpenStudy (anonymous):

\[b \times b=144\]

OpenStudy (anonymous):

there's no dividing here. it's squares and square roots.

OpenStudy (anonymous):

what times itself is 144?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

:-) !!!!!!!!!!!

OpenStudy (anonymous):

so it's a 3-4-5 triangle with sides 4x as large: 12-16-20

OpenStudy (anonymous):

Thank you was confuse with dividing instad of square root

OpenStudy (anonymous):

no problem.

OpenStudy (anonymous):

I have one more question

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

Square root ^x15would that be 3

OpenStudy (anonymous):

complete the expression

OpenStudy (anonymous):

I can't understand what you wrote...

OpenStudy (anonymous):

what does "square root ^x15" mean?

OpenStudy (anonymous):

its a square root sign 5x on the outside andx^15 inside

OpenStudy (anonymous):

\[\sqrt[5]{x^{15}}\]???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\sqrt[5]{x^{15}}=(x^{15})^\frac{1}{5}=x^{15\times\frac{1}{5}}=x^3\]

OpenStudy (anonymous):

So i just left off the x